A sample of water was alkaline both to phenolphthalein and methyl orange. 100 mL of this water sample required 20 mL of N/50 HCl for phenolphthalein end point and another 15 mL for complete neutralization. Elucidate the type and strength of alkalinity present in the water sample


1
Expert's answer
2021-06-02T06:04:23-0400

Here,diacidic base is present which gives two end points one with phenolphthalein and another with methyl orange.

Now, for phenolphthalein end point

"\frac{1}{2}M" equivalent to diacidic base"=1M" equivalent of 15mL"\frac{N}{50}HCl"

or "\frac{1}{2}M" equivalent of diacidic base=


"\frac{1×20}{50×1000}=\frac{1}{2500}M equivalent"


or, 1M equivalent of diacidic base="2×\frac{1}{2500}M equivalent of HCl"


But,

For Methyl Orange end point

1M equivalent of monoacidic base formed=1M equivalent of 15mL"\frac{N}{50}HCl"

or 1M equivalent of monoacidic base formed

"=\frac{1×15}{50×1000}=\frac{3}{10,000}M of HCl"


Thus ,normality of diacidic alkalinity present in 100mL sample of water is given by :-

"\frac{2×1000}{2500×100}=\frac{2}{250}=\frac{1}{125}N or 0.008N"